OpenStudy (anonymous):

@Saifoo.Khan, how did u get the answer for "sin ϴ= 3/7. find csc ϴ."??

OpenStudy (anonymous):

csc is a inverse

OpenStudy (anonymous):

of sine

OpenStudy (anonymous):

Csc(x)=1/Sin(x) So you have: 1/3/7=7/3

OpenStudy (saifoo.khan):

just inverse it.

OpenStudy (anonymous):

yup..

OpenStudy (anonymous):

Oh i c that now :) Thanks!

OpenStudy (saifoo.khan):

this might help

OpenStudy (anonymous):

yeah, it did, ty

OpenStudy (saifoo.khan):

uw

OpenStudy (saifoo.khan):

\begin{array}l\color{#624664}{\text{=}}\color{#644519}{\text{]}}\color{#62a251}{\text{}}\end{array}

OpenStudy (saifoo.khan):

hi hero!

OpenStudy (anonymous):

saifoo khan, always coming when the females show up...lol

OpenStudy (saifoo.khan):

haha, thats wht u think..

OpenStudy (anonymous):

School must be over on a Sunday....lol All the females and Saifoo show up after 5pm, lol

OpenStudy (saifoo.khan):

haha, its monday here! & i've already finishd my School. :P Im on vacations!